OpenStudy (anonymous):

What is the solution to the equation 9(x + 1) = 27 ? x = 2.5 x = 0.5 x = –0.5 x = –3.5

OpenStudy (akashdeepdeb):

Open the brackets! 9x + 9 = 27!! Now take the 9 to the RHS! And then divide the RHS by 9! Get the answer?

OpenStudy (anonymous):

Oops i typed that wrong. 9^(x + 1) = 27 Is it still the same process? @AkashdeepDeb

OpenStudy (akashdeepdeb):

No now you know that 9^3 = 27 right? So 9^(x+1) = 9^3 Now as the bases are same equate the powers! Get your answer! :)

OpenStudy (anonymous):

9^3 = 729

OpenStudy (anonymous):

Express them in powers of 3

OpenStudy (akashdeepdeb):

How much is 9? It is (3^2)^3 = 3^6 ! :)

OpenStudy (anonymous):

9^a = 3^(2a)